OpenStudy (anonymous):

What quantity of gas, in moles, is contained in 2.21L at STP?

OpenStudy (anonymous):

P=10^5 pa T= 25+273 K n=1 and R=8.314 m3pa.mol-1.K-1

OpenStudy (anonymous):

around 0.1 mol

OpenStudy (anonymous):

P=10^5 pa T= 25+273 K R=8.314 m3pa.mol-1.K-1 AND V=2.21*10^-3 m3 SO n=0.1

OpenStudy (anonymous):

PV=nRT